# Nathalie Gimenes PUC-Rio Tonghui Qi

### v1

@@ 3.1 Assumptions for AQR estimation
(i) The auction variables $\left\{\left(B_{l},X_{l}\right)\right\}_{l=1}^{L}$ are a random sample. For some $F\left(\cdot|X\right)$ and $G\left(\cdot|X\right)$ that satisfy Assumptions M1 and M2(i), $B_{l}$ takes value in $V$ and have conditional CDF $G\left(\cdot|X\right)$. $X_{l}$ takes value in $X\subseteq R^{D}$ such that $X$ is compact, and the eigenvalues of $E\left[X_{l}X_{l}^{\top}\right]$ are bounded away from zero and infinity.
@@ 3.2 Estimator of quantile function
where $\rho_{\alpha}\left(t\right)=t\left(\alpha-\mathbf{1}\left[t<0\right]\right)$ is the check function. The minimizer of the sample counterpart to the expectation above is the classic quantile regression estimator of Koenker and Bassett, (1978). This estimator is known to not perform well at extreme quantiles ($\alpha$ close to 0 or $1$). Moreover, it is piecewise linear and different estimators for quantile derivatives are required that complicates analysis of statistics that involve quantile level and its derivatives. Instead, let us consider $B\left(\cdot|X\right)$ over $\left[\alpha-h,\alpha+h\right]\cap\left[0,1\right]$, $\left\{B\left(\tau|X\right),\tau\in\left[\alpha-h,\alpha+h\right]\cap\left[0,1\right]\right\}$, which minimizes
@@
over any functions $q\left(\cdot,X\right)$ since $K\left(\cdot\right)$ is non-negative. In the same spirit as a local polynomial estimator (e.g., Fan and Gijbels, (1996)), the AQR approach, motivated by the Taylor’s expansion, estimates the quantile coefficients and their derivatives simultaneously. Specifically, with $B\left(\alpha+th|x\right)=\sum\limits_{j=0}^{s}x_{1}^{\top}\beta^{\left(j\right)}\left(\alpha\right)\frac{\left(th\right)^{j}}{j!}+O\left(h^{s+1}\right)$ in mind, consider the following objective function,
@@
where $P\left(x,t\right)=\pi\left(t\right)\otimes x_{1}$ with $\pi\left(t\right)=\left[1,t,\ldots,\frac{t^{s}}{s!}\right]^{\top}$ and $x_{1}=\left[1,x^{\top}\right]^{\top}$. [^5] It is worth pointing out a subtle difference between our objective function relative to GG22 here. Despite of us both assuming $\left(s+1\right)$-times continuous differentiability of $V\left(\cdot\right)$, we make a polynomial approximation up to the $s$-th power term while GG22 goes up to the $\left(s+1\right)$-th power term. This is because the quantile function of the optimal first price auction bid has one more derivative than the quantile function of the private value – see equation (2.4) in GG22, which is given in the previous footnote.
@@
Before giving the statistical properties of $\widehat{V}\left(\alpha|x\right)$, we provide pointwise properties on $\widehat{B}\left(\alpha|x\right)$. This may be of independent interest for the AQR estimator, noting that these are not given in GG22.
@@
The appendix gives expressions for the AQR estimator decomposition, $\left(J_{B}\left(\alpha,x\right),J_{S}\left(\alpha,x\right),J_{R}\left(\alpha,x\right)\right)$, in equations (18) to (20). These terms respectively represent the bias, leading stochastic term, and remainder term of $\widehat{B}\left(\alpha|x\right)-B\left(\alpha|x\right)$. Note that the limiting distribution of $\sqrt{L}J_{S}\left(\alpha,x\right)$ is the same as the standard quantile regression estimator’s without smoothing (for example, see Chapter 4 of Koenker, (2005)), so that the AQR estimator has the same first order asymptotic property as the Koenker and Bassett’s estimator when $Lh^{2\left(s+1\right)}=o\left(1\right)$.
@@ 3.3 Estimators of related functions
Using the above equation for estimation requires estimators for $\left(V^{\left(1\right)}\left(\cdot\right),V^{-1}\left(\cdot\right)\right)$. Since the convergence rates for an estimator of $V^{\left(j\right)}\left(\cdot\right)$ may be generally useful for semiparametric estimation, we provide convergence rates for these objects as well. In what follows, we provide the relations between $\left(V^{\left(j\right)}\left(\cdot\right),V^{-1}\left(\cdot\right)\right)$ and $\left(B^{\left(j\right)}\left(\cdot\right),B^{-1}\left(\cdot\right)\right)$. Then, we define our estimators as transformations of the AQR estimators of $\left(B^{\left(j\right)}\left(\cdot\right),B^{-1}\left(\cdot\right)\right)$ and give their rates of convergence. As done for the quantile function previously, it will be instructive to first provide some statistical properties of $\left(\widehat{B}^{\left(j\right)}\left(\cdot\right),\widehat{B}^{-1}\left(\cdot\right)\right)$ as lemmas.
@@ 3.3.1 Derivatives of the quantile function
which hold a.s. for all $\alpha\in\left[0,1\right]$, it gets cumbersome quickly for higher derivatives. Since we are only providing the uniform convergence rates rather than pointwise properties of $V^{\left(j\right)}(\cdot)$, it suffices to know that:
@@
Since the AQR approach readily gives estimator for $B^{\left(j\right)}(\alpha|x)$ for $j=1,\ldots,s$, we estimate $V^{\left(j\right)}(\alpha|x)$ by
@@
with $\mathsf{S}_{j}=S_{j}\otimes\mathrm{I}_{D+1}$ and $S_{j}$ is a row vector of size $\left(s+1\right)$ consists of 0’s in every component other than $1$ in its $\left(j+1\right)$-th entry.
Lemma 2 gives the convergence rate for the AQR estimator of $\widehat{B}^{\left(j\right)}\left(\cdot\right)$.
@@
Notice that Lemma 2 imposes the same bandwidth condition as GG22, which we alluded earlier. Indeed, GG22 has given the same convergence rate as the above when $j=1$, see equation (5.7) in their Theorem 2.777The order of their bias is written as $h^{s+1}$, which is a result of their bid’s quantile function having one more derivative than ours. The component of the convergence rate that corresponds to the stochastic term is $\sqrt{\frac{\log L}{Lh^{2j-1}}}$, which coincides with the usual rates of the $\left(j-1\right)$-th derivative of a kernel density estimator; this finding is reassuring given the identity between the density and derivative of the quantile. The worsening of the bias rate with higher derivatives also mirrors standard local polynomial estimators. Since $\widehat{V}^{\left(j\right)}\left(\cdot\right)$ is a smooth in $\left\{\widehat{B}^{\left(k\right)}\left(\cdot\right)\right\}_{k=1}^{j}$, its rate of convergence then follows that of $\widehat{B}^{\left(j\right)}\left(\cdot\right)$.
@@ 3.3.2 Inverse of the quantile function
The inverse of the quantile function is the CDF. The relation between the inverse of private value and winning bid quantiles are given by the inverse of composite functions formula applied to (1):
@@
First, $\widehat{B}^{-1}\left(t|x\right)$ can be studied as the inverse of $\widehat{B}\left(t|x\right)$, through the map $\Psi:\ell^{\infty}\left(\left[0,1\right]\right)\mapsto\ell^{\infty}\left(\mathcal{V}\right)$, such that $\Psi\left(\pi\right)\left(t\right)=\inf\left\{a\in\left[0,1\right]:\pi\left(a\right)\geq t\right\}$ for $t\in\mathcal{V}$ and $\pi\left(\cdot\right)\in\ell^{\infty}\left(\left[0,1\right]\right)$. Here we use $\ell^{\infty}\left(\mathcal{A}\right)$ to denote the space of bounded functions on $\mathcal{A}\subseteq\mathbb{R}$. It is well known that the Hadamard derivative of $\Psi$ exists and linearization methods apply (van der Vaart and Wellner, (2023), Lemma 3.10.21). Particularly, the leading term from linearizing $\widehat{B}^{-1}\left(t|x\right)-B^{-1}\left(t|x\right)$ is,
@@
Lemma 3 gives, along side the uniform convergence rate, the pointwise properties of $\widehat{B}^{-1}\left(t|x\right)$ for $t\in int\left(\mathcal{V}\right)$, denoting in the interior of $\mathcal{V}$. The latter may be of independent interest, as $\widehat{B}^{-1}\left(t|x\right)$ is a flexible yet low-dimensional general estimator for the conditional CDF.
@@
$\phi\left(a\right)$ $=$ $Ia^{I-1}-\left(I-1\right)a^{I}$
$\phi^{\prime}\left(a\right)$ $=$ $I\left(I-1\right)a^{I-2}\left(1-I\left(I-1\right)a\right)$
Second, we apply $\phi^{-1}\left(\cdot\right)$ to $\widehat{B}^{-1}\left(\cdot\right)$. There is a potential complication when deriving uniform convergence in this step as $\phi^{-1}\left(\cdot\right)$ is continuously differentiable on $\left(0,1\right)$ but not at the boundaries. This can be seen from inspecting the derivative of $\phi^{-1}\left(\cdot\right)$, which is $1/\phi^{\left(1\right)}\left(\phi^{-1}\left(\cdot\right)\right)$, as we have $\phi^{\left(1\right)}\left(0\right)=0$ when $I>2$ and $\phi^{\left(1\right)}\left(1\right)$ is 0 for all $I$. It should be noted too that the bias and variance of $\widehat{B}^{-1}\left(\cdot\right)$ go to zero as $t\rightarrow\underline{v}$, and also $t\rightarrow\overline{v}$ for the variance. These faster convergence rates may mitigate potential irregularity issues for the de-meaned component of $\widehat{V}^{-1}\left(\cdot\right)$ as well as the lower boundary bias. Nevertheless, we do not need these aspects to derive the large sample properties of our risk aversion parameter in the next section, and a comprehensive study on uniform properties of such transformed AQR estimator is beyond the scope of this paper.
The next proposition gives the uniform convergence rate for $\widehat{V}^{-1}\left(\cdot\right)$ on an inner interval of $\mathcal{V}$, denoted by $\mathcal{V}_{\delta}$ and defined as $\left[\underline{v}+\delta,\overline{v}-\delta\right]$ for $\delta\in\left(0,\left(\overline{v}-\underline{v}\right)/2\right)$.
@@
4 Risk aversion estimator
We now assume the utility function takes a parametric form: $\theta\mapsto U_{\theta}\left(\cdot\right)$, for some $\theta\in\Theta\subset\mathbb{R}$ that represents an Arrow-Pratt measure of risk aversion. We can then construct an objective function for estimating the risk parameter from the first-order condition in (9). To do this, let us use $\mathcal{D}_{1}\left(\mathcal{A_{\delta}}\times\mathcal{X}\right)$ and $\mathcal{D}_{2}\left(\mathcal{V_{\delta}}\times\mathcal{X}\right)$ to denote classes of functions whose images are $\mathcal{V}_{\delta}$ and $\mathcal{A}_{\delta}$ respectively. We denote candidates for the derivative and inverse of the conditional quantile function of $V_{l}$ given $X_{l}$ by $\psi_{1}\left(\cdot\right)\in\mathcal{D}_{1}\left(\mathcal{A_{\delta}}\times\mathcal{X}\right)$ and $\psi_{2}\left(\cdot\right)\in\mathcal{D}_{2}\left(\mathcal{V_{\delta}}\times\mathcal{X}\right)$ respectively. Here, we use $\mathcal{V}_{\delta}=\left[\underline{v}+\delta,\overline{v}-\delta\right]$ and $\mathcal{A}_{\delta}=\left[\delta,1-\delta\right]$ for small $\delta$. We restrict the support of the quantile and valuation for the reason discussed at the end of Section 3. Note that $\delta$ in $\mathcal{V}_{\delta}$ and $\mathcal{A}_{\delta}$ can generally be different. Moreover, the supports of (bidder’s and seller’s) valuation (and the reserve price) can depend on $X$. Incorperating these is conceptually straightforward. We forego the more general notations for simplicity of presentation.
@@ 4 Risk aversion estimator
Our structural assumption will, by design, require that the first order condition in (9) coincides with $q\left(z,\theta_{0},\psi_{0}\right)$ for some $\left(\theta_{0},\psi_{0}\right)$ for all $z$ that is consistent with the auction model describe in Section 2. This is suggestive for a least squares or minimum distance type objective function for estimating $\theta_{0}$.
@@
Given our usage of empirical process methods to proving the asymptotic results, we write $Q$ as an integral, making clear that only $Z$ is being integrated out, so that $Q\left(\theta,\psi\right)$ is a random variable if either $\theta\$or $\psi$ is random. We add that, more precisely, $P_{Z}$ can be understood as a conditional distribution for $R\in\mathcal{V}_{\delta}$ for the purpose of the proofs, although there is no data trucation in practice. Related to the latter point, we emphasize that, we are not at risk of identification loss with our minimum distance approach by working on $\left(\mathcal{A}_{\delta},\mathcal{V}_{\delta}\right)$ instead of $\left(\left[0,1\right],\mathcal{V}\right)$, which should be contrasted with choosing moments in a conditional moment model (e.g., see Domínguez and Lobato, (2004)), as a single value of $z\in\mathcal{Z}$ identifies $\theta_{0}$ via (13).
Our estimation problem here is a semiparametric one, as we are interested in the finite dimensional parameter $\theta_{0}$ in the presence of infinite dimensional nuisance functions. We denote the estimator for the latter by $\widehat{\psi}$, which consists of $\widehat{\psi}_{1}=\widehat{V}^{\left(1\right)}$ and $\widehat{\psi}_{2}=\widehat{V}^{-1}$, respectively defined as in 3.3.1 and 3.3.2 using a kernel function that satisfies Assumption Q(iii). We then define our estimator for $\theta_{0}$ to be any $\widehat{\theta}$ that satisfies:
$\left|Q_{L}\left(\widehat{\theta},\widehat{\psi}\right)\right|\leq\inf_{\theta\in\Theta}\left|Q_{L}\left(\theta,\widehat{\psi}\right)\right|+o_{p}\left(1/\sqrt{L}\right).$ (14)
@@
Assumption S1 consists of structural assumptions and regularity conditions on the variables in the model.
The condition $q\left(Z,\theta_{0},\psi_{0}\right)=0$ in S1(i) assumes correct model specification and, together with S1(ii), they imply the distribution of the data can be rationalized by the ascending auction model described in Section 2. Particularly, it implies that $\psi_{10}\left(\alpha,x\right)=V^{\left(1\right)}\left(\alpha|x\right)=\phi^{\left(1\right)}\left(\alpha\right)B^{\left(1\right)}\left(\phi\left(\alpha\right)|x\right)$ for $\left(\alpha,x\right)\in\mathcal{A}_{\delta}\times\mathcal{X}$ and $\psi_{20}\left(t,x\right)=V^{-1}\left(t|x\right)=\phi^{-1}\left(t\right)B^{-1}\left(\phi\left(t\right)|x\right)$ for $\left(t,x\right)\in\mathcal{V}_{\delta}\times\mathcal{X}$.
S1(iii) imposes a parametric assumption on the utility function that satisfies M2(iv). Part (a) gives an interpretation for $\theta$ to be an Arrow–Pratt coefficient where higher $\theta$ means higher degree of risk aversion in the Arrow–Pratt sense as described in Section 2.2. For example, the risk parameters in CARA and CRRA utility functions satisfy this condition (Ross, (1981)). Part (b) is a regularity condition requiring uniform squared integrability. In our application we use the CRRA utility function, which is defined in (6). It should be noted that such utility function is defined for all $\theta\in\mathbb{R}$, where $\theta>0$, $\theta=0,$ and $\theta<0$ represent risk-averse, risk-neutral, and risk-loving preferences respectively. Under the CRRA preference, S1(iii) requires $\theta\geq 0$. The sole purpose of this is to ensure we can apply Proposition 2, which we use to prove Lemma 4 that shows $\theta_{0}$ is identified as the minimizer of $Q$ under primitive conditions. Our estimation procedure does not restrict the parameter space to be non-negative. Importantly, if the implication of Lemma 4 is assumed to hold as a high-level identification condition, the asymptotic theory for our estimator applies for $\theta_{0}<0$ without any modification. [^8] Our simulation study does not suggest any identification issue when $\theta_{0}<0$, and the estimator in the risk-loving case behaves in the same was as the risk-neutral and risk-averse cases qualitatively. These additional simulation results are available upon request. We can therefore abstract away from the inference issues that arise when parameter is on the boundary such as those discussed in Andrews, (2001). Lastly, on parametric modelling, we can also embed observables in the risk measure, for example, by replacing $\theta$ with a linear index of the seller’s characteristics.
@@
where $c_{Y_{1}}$ and $c_{Y_{2}}$ are functions of $Z$, and $\circ$ denotes the composition of functions. See (28) and (29) in Appendix C respectively for the explicit forms of $c_{Y_{1}}$ and $c_{Y_{2}}$.
@@
It is worth noting that our bandwidth requirement of $\log^{2}L=o\left(Lh^{3}\right)$ coincides with the condition in Theorem 4 of GG22, which establishes asymptotic normality for a similar integral functional that contains the bid’s quantile function and its derivative. This is reassuring, as they estimate their functional using AQR estimators that have the same convergence rates as $\widehat{\psi}_{20}$ and $\widehat{\psi}_{10}$Ḟor example, when the bandwidth decays at the rate $\log L^{1/2}L^{-\varsigma}$ for some $\varsigma>0$, Lemma 6 requires $\frac{1}{4s}<\varsigma<\frac{1}{3}$ for it to hold.
@@
In practice, we recommend that inference on $\theta_{0}$ be performed using the ordinary nonparametric bootstrap. Chen et al., (2003) provide high-level conditions under which the asymptotic distribution of a two-step semiparametric estimator can be consistently bootstrapped. We conjecture that our setup is amenable to such a result. Indeed, we apply this in our Monte Carlo study and find that bootstrap standard errors for $\widehat{\theta}$ perform very well. A formal proof of bootstrap validity when AQR estimators are used to estimate nuisance functions in a general two-step semiparametric procedure is, however, beyond the scope of this paper.

### v2

@@ 3.1 Assumptions for AQR estimation
(i) The auction variables $\left\{\left(B_{l},X_{l}\right)\right\}_{l=1}^{L}$ are a random sample. For some $F\left(\cdot|X\right)$ and $G\left(\cdot|X\right)$ that satisfy Assumptions M1 and M2(i), $B_{l}$ takes value in $\mathcal{V}$ and have conditional CDF $G\left(\cdot|X\right)$. $X_{l}$ takes value in $\mathcal{X}\subseteq R^{D}$ such that $\mathcal{X}$ is compact, and the eigenvalues of $E\left[X_{l}X_{l}^{\top}\right]$ are bounded away from zero and infinity.
@@ 3.2 Estimator of quantile function
where $\rho_{\alpha}\left(t\right)=t\left(\alpha-\mathbf{1}\left[t<0\right]\right)$ is the check function. The minimizer of the sample counterpart to the expectation above is the classic quantile regression estimator of Koenker and Bassett, (1978). This estimator is known to not perform well when $\alpha$ is close to 0 or $1$. Moreover, it is piecewise linear (in $\alpha$) and different estimators for quantile derivatives are required that complicates analysis of statistics that involve both quantile level and its derivatives estimates.
Instead, let us consider $B\left(\cdot|X\right)$ over $\left[\alpha-h,\alpha+h\right]\cap\left[0,1\right]$, $\left\{B\left(\tau|X\right),\tau\in\left[\alpha-h,\alpha+h\right]\cap\left[0,1\right]\right\}$, which minimizes
@@
over any functions $q\left(\cdot,X\right)$ since $K\left(\cdot\right)$ is non-negative. In the same spirit as a local polynomial estimator (e.g., Fan and Gijbels, (1996)), the AQR approach estimates the quantile coefficients and their derivatives simultaneously. Specifically, with $B\left(\alpha+th|x\right)=\sum\limits_{j=0}^{s}x_{1}^{\top}\beta^{\left(j\right)}\left(\alpha\right)\frac{\left(th\right)^{j}}{j!}+O\left(h^{s+1}\right)$ in mind, consider the following objective function,
@@
where $b\in\mathbb{R}^{\left(s+1\right)\left(D+1\right)}$ and $P\left(x,t\right)=\pi\left(t\right)\otimes x_{1}$ with $\pi\left(t\right)=\left[1,t,\ldots,\frac{t^{s}}{s!}\right]^{\top}$ and $x_{1}=\left[1,x^{\top}\right]^{\top}$. [^6] There is a subtle difference between our objective function relative to GG22 here. Despite of us both assuming $\left(s+1\right)$-times continuous differentiability of $V\left(\cdot\right)$, we make a polynomial approximation up to the $s$-th power term while GG22 goes up to the $\left(s+1\right)$-th power term. This is because the quantile function of the optimal first price auction bid has one more derivative than the quantile function of the private value—see equation (2.4) in GG22, which is given in the previous footnote.
@@
Since $\widehat{V}\left(\alpha|x\right)$ inherits the properties of $\widehat{B}\left(\alpha|x\right)$, we first provide the statistical properties of the latter as a lemma. Moreover, Lemma 1 contains pointwise properties for an AQR estimator, which is not given in GG22, as they only provide the uniform convergence rate—see Theorem D.1 in their paper.
@@
The Appendix gives expressions for $\left(J_{B}\left(\alpha,x\right),J_{S}\left(\alpha,x\right),J_{R}\left(\alpha,x\right)\right)$ in equations (17) to (19). These terms respectively represent the bias, leading stochastic term, and remainder term of $\widehat{B}\left(\alpha|x\right)-B\left(\alpha|x\right)$. Note that the limiting distribution of $\sqrt{L}J_{S}\left(\alpha,x\right)$ is the same as the standard quantile regression estimator’s without smoothing (for example, see Chapter 4 of Koenker, (2005)), so that the AQR estimator has the same first order asymptotic property as the Koenker and Bassett’s estimator when $Lh^{2\left(s+1\right)}=o\left(1\right)$.
@@ 3.3 Estimators of related functions
Using the above equation for estimation requires estimators for $\left(V^{\left(1\right)}\left(\cdot\right),V^{-1}\left(\cdot\right)\right)$. Since the convergence rates for an estimator of $V^{\left(j\right)}\left(\cdot\right)$ is useful for semiparametric estimation, we begin by providing convergence rates for them. In what follows, we provide the relations between $\left(V^{\left(j\right)}\left(\cdot\right),V^{-1}\left(\cdot\right)\right)$ and $\left(B^{\left(j\right)}\left(\cdot\right),B^{-1}\left(\cdot\right)\right)$. Then, we define our estimators as transformations of the AQR estimators of $\left(B^{\left(j\right)}\left(\cdot\right),B^{-1}\left(\cdot\right)\right)$ and give their rates of convergence.
@@ 3.3.1 Derivatives of the quantile function
which hold a.s. for all $\alpha\in\left[0,1\right]$, they get cumbersome quickly for higher derivatives. Since we are only interested in the uniform convergence rates rather than pointwise properties here, it therefore suffices to know that:
@@
The AQR approach readily estimates $B^{\left(j\right)}(\alpha|x)$ for $j=1,\ldots,s$, which gives
@@
with $\mathsf{S}_{j}=S_{j}\otimes\mathrm{I}_{D+1}$ and $S_{j}$ is a row vector of size $\left(s+1\right)$ consists of 0’s in every component other than $1$ in its $\left(j+1\right)$-th entry. Lemma 2 gives the convergence rate for the AQR estimator of $\widehat{B}^{\left(j\right)}\left(\cdot\right)$.
@@
Notice that Lemma 2 imposes the same bandwidth condition as GG22, which we alluded earlier. Indeed, GG22 has given the same convergence rate as the above when $j=1$, see equation (5.7) in their Theorem 2.888The order of their bias is written as $h^{s+1}$, which is a result of their bid’s quantile function having one more derivative than ours. The component of the convergence rate that corresponds to the stochastic term is $\sqrt{\frac{\log L}{Lh^{2j-1}}}$, which coincides with the usual rates of the $\left(j-1\right)$-th derivative of a kernel density estimator; this finding is reassuring given the identity between the density and derivative of the quantile. The worsening of the bias rate with higher derivatives also mirrors standard local polynomial estimators. Since $\widehat{V}^{\left(j\right)}\left(\cdot\right)$ is a smooth mapping from $\left\{\widehat{B}^{\left(k\right)}\left(\cdot\right)\right\}_{k=1}^{j}$, its rate of convergence then follows that of $\widehat{B}^{\left(j\right)}\left(\cdot\right)$.
@@ 3.3.2 Inverse of the quantile function
The relation between the inverse of private value and winning bid quantiles is given by the inverse of composite functions formula applied to (1):
@@
First, $\widehat{B}^{-1}\left(t|x\right)$ can be studied as the inverse of $\widehat{B}\left(t|x\right)$ through the map $\Psi:\ell^{\infty}\left(\left[0,1\right]\right)\mapsto\ell^{\infty}\left(\mathcal{V}\right)$, such that $\Psi\left(\pi\right)\left(t\right)=\inf\left\{a\in\left[0,1\right]:\pi\left(a\right)\geq t\right\}$ for $t\in\mathcal{V}$ and $\pi\left(\cdot\right)\in\ell^{\infty}\left(\left[0,1\right]\right)$. We use $\ell^{\infty}\left(\mathcal{A}\right)$ to denote the space of bounded functions on $\mathcal{A}\subseteq\mathbb{R}$. It can be shown that the Hadamard derivative of $\Psi$ exists and linearization methods apply (van der Vaart and Wellner, (2023), Lemma 3.10.21). Particularly, the leading term in $\widehat{B}^{-1}\left(t|x\right)-B^{-1}\left(t|x\right)$ is,
@@
Lemma 3 gives the statistical properties of $\widehat{B}^{-1}\left(t|x\right)$ for $t$ in the interior of $\mathcal{V}$, denoted by $int\left(\mathcal{V}\right)$. This result may be of independent interest, as $\widehat{B}^{-1}\left(t|x\right)$ is a flexible yet low-dimensional general estimator for the conditional CDF.
@@
Second, we apply $\phi^{-1}\left(\cdot\right)$ to $\widehat{B}^{-1}\left(\cdot\right)$. There is a potential complication when deriving uniform convergence in this step, as $\phi^{-1}\left(\cdot\right)$ is continuously differentiable on $\left(0,1\right)$ but not at the boundaries. This can be seen from inspecting the derivative of $\phi^{-1}\left(\cdot\right)$, which is $1/\phi^{\left(1\right)}\left(\phi^{-1}\left(\cdot\right)\right)$, as we have $\phi^{\left(1\right)}\left(0\right)=0$ when $I>2$ and $\phi^{\left(1\right)}\left(1\right)$ is 0 for all $I$. It should be noted too that the bias and variance of $\widehat{B}^{-1}\left(\cdot\right)$ go to zero as $t\rightarrow\underline{v}$, and also $t\rightarrow\overline{v}$ for the variance. These faster convergence rates may mitigate potential irregularity issues for the de-meaned component of $\widehat{V}^{-1}\left(\cdot\right)$ as well as the lower boundary bias. Nevertheless, we do not need these aspects to derive the large sample properties of our risk aversion parameter in the next section, and a comprehensive study on uniform properties of such transformed AQR estimator is beyond the scope of this paper.
The next proposition gives the uniform convergence rate for $\widehat{V}^{-1}\left(\cdot\right)$ on an inner interval of $\mathcal{V}$, denoted by $\mathcal{V}_{\delta}$ that is defined as $\left[\underline{v}+\delta,\overline{v}-\delta\right]$ for $\delta\in\left(0,\left(\overline{v}-\underline{v}\right)/2\right)$.
@@
4 Risk-aversion estimator
We now assume the utility function takes a parametric form: $\theta\mapsto U_{\theta}\left(\cdot\right)$, for some $\theta\in\Theta\subset\mathbb{R}$ that represents an Arrow-Pratt measure of risk aversion. We can then construct an objective function for estimating the risk parameter from the first-order condition in (9). To do this, let us use $\mathcal{D}_{1}\left(\mathcal{A_{\delta}}\times\mathcal{X}\right)$ and $\mathcal{D}_{2}\left(\mathcal{V_{\delta}}\times\mathcal{X}\right)$ to denote classes of functions whose images are $\mathcal{V}_{\delta}$ and $\mathcal{A}_{\delta}$ respectively. We denote candidates for the derivative and inverse of the conditional quantile function of $V_{l}$ given $X_{l}$ by $\psi_{1}\left(\cdot\right)\in\mathcal{D}_{1}\left(\mathcal{A_{\delta}}\times\mathcal{X}\right)$ and $\psi_{2}\left(\cdot\right)\in\mathcal{D}_{2}\left(\mathcal{V_{\delta}}\times\mathcal{X}\right)$ respectively. Here, we use $\mathcal{V}_{\delta}=\left[\underline{v}+\delta,\overline{v}-\delta\right]$ and $\mathcal{A}_{\delta}=\left[\delta,1-\delta\right]$ for small $\delta$. We restrict the support of the quantile and valuation for the reasons discussed at the end of Section 3. Note that $\delta$ in $\mathcal{V}_{\delta}$ and $\mathcal{A}_{\delta}$ can generally be different. Moreover, the supports of (bidder’s and seller’s) valuation (and the reserve price) can depend on $X$. Incorperating these is conceptually straightforward. We forego the more general notations for simplicity of presentation.
@@ 4 Risk-aversion estimator
Our structural assumption requires that the first order condition in (9) coincides with $q\left(z,\theta_{0},\psi_{0}\right)$ for some $\left(\theta_{0},\psi_{0}\right)$ for all $z$ that is consistent with the auction model describe in Section 2. This is suggestive for a minimum distance type objective function for estimating $\theta_{0}$.
@@
Given our usage of empirical process methods to proving the asymptotic results, we write $Q$ as an integral, making clear that only $Z$ is being integrated out, so that $Q\left(\theta,\psi\right)$ is a random variable if either $\theta\$or $\psi$ is random. We add that, more precisely, $P_{Z}$ can be understood as a conditional distribution for $R\in\mathcal{V}_{\delta}$ for the purpose of the proofs, although there is no data truncation in practice. Related to the latter point, we emphasize that, we are not at risk of identification loss with our minimum distance approach by working on $\left(\mathcal{A}_{\delta},\mathcal{V}_{\delta}\right)$ instead of $\left(\left[0,1\right],\mathcal{V}\right)$, which should be contrasted with choosing moments in a conditional moment model (e.g., see Domínguez and Lobato, (2004)), as a single value of $z\in\mathcal{Z}$ identifies $\theta_{0}$ via (13).
[…]
@@
Assumption S1 consists of structural assumptions and regularity conditions on the variables in the model. The condition $q\left(Z,\theta_{0},\psi_{0}\right)=0$ in S1(i) assumes correct model specification and, together with S1(ii), they imply the distribution of the data can be rationalized by the ascending auction model described in Section 2. Particularly, it implies that $\psi_{10}\left(\alpha,x\right)=V^{\left(1\right)}\left(\alpha|x\right)=\phi^{\left(1\right)}\left(\alpha\right)B^{\left(1\right)}\left(\phi\left(\alpha\right)|x\right)$ for $\left(\alpha,x\right)\in\mathcal{A}_{\delta}\times\mathcal{X}$ and $\psi_{20}\left(t,x\right)=V^{-1}\left(t|x\right)=\phi^{-1}\left(t\right)B^{-1}\left(\phi\left(t\right)|x\right)$ for $\left(t,x\right)\in\mathcal{V}_{\delta}\times\mathcal{X}$.
S1(iii) imposes a parametric assumption on the utility function that satisfies M2(iv). Part (a) gives an interpretation for $\theta$ to be an Arrow–Pratt coefficient where higher $\theta$ means higher degree of risk aversion in the Arrow–Pratt sense as described in Section 2.2. For example, the risk parameters in CARA and CRRA utility functions satisfy this condition (Ross, (1981)). Part (b) is a regularity condition requiring uniform squared integrability. In our application we use the CRRA utility function, which is defined in (6). It should be noted that such utility function is defined for all $\theta\in\mathbb{R}$, where $\theta>0$, $\theta=0,$ and $\theta<0$ represent risk-averse, risk-neutral, and risk-loving preferences respectively. Under the CRRA preference, S1(iii) requires $\theta\geq 0$. The sole purpose of this is to ensure we can apply Proposition 2, which we use to prove Lemma 4 that shows $\theta_{0}$ is identified as the minimizer of $Q$ under primitive conditions. Our estimation procedure does not restrict the parameter space to be non-negative. Importantly, suppose the implication of Lemma 4 holds as a high-level identification condition, the asymptotic theory for our estimator applies for $\theta_{0}<0$ without any modification. [^9] Our simulation study does not suggest any identification issue when $\theta_{0}<0$, and the estimator in the risk-loving case behaves in the same was as the risk-neutral and risk-averse cases qualitatively. These additional simulation results are available upon request. We can therefore abstract away from the inference issues that arise when parameter is on the boundary such as those discussed in Andrews, (2001). Note also that the parametric model can be enriched by embedding observables in the risk measure, for example, by replacing $\theta$ with a linear index of the seller’s or auction characteristics.
@@
where $c_{Y_{1}}$ and $c_{Y_{2}}$ are functions of $Z$, and $\circ$ denotes the composition of functions. See equations (27) and (28) in the Appendix respectively for the explicit forms of $c_{Y_{1}}$ and $c_{Y_{2}}$.
@@
It is worth noting that our bandwidth requirement of $\log^{2}L=o\left(Lh^{3}\right)$ coincides with the condition in Theorem 4 of GG22, which establishes asymptotic normality for a similar integral functional that contains the bid’s quantile function and its derivative. This is reassuring, as they estimate their functional using AQR estimators that have the same convergence rates as $\widehat{\psi}_{20}$ and $\widehat{\psi}_{10}$. For example, when the bandwidth decays at the rate $\log L^{1/2}L^{-\varsigma}$ for some $\varsigma>0$, Lemma 6 requires $\frac{1}{4s}<\varsigma<\frac{1}{3}$ for it to hold.
@@
In practice, inference on $\theta_{0}$ be performed using a nonparametric bootstrap. Chen et al., (2003) provide high-level conditions under which the asymptotic distribution of a two-step semiparametric estimator can be consistently bootstrapped. We conjecture that our setup is amenable to such a result. Indeed, we apply this in our Monte Carlo study and find that bootstrap standard errors for $\widehat{\theta}$ perform very well. A formal proof of bootstrap validity when AQR estimators are used to estimate nuisance functions in a general two-step semiparametric procedure is, however, beyond the scope of this paper.